NASA is conducting an experiment to find out the fraction of people who black out at G forces greater than 6.

Step 2 of 2 :  

Suppose a sample of 977 people is drawn. Of these people, 439 passed out. Using the data, construct the 99% confidence interval for the population proportion of people who black out at G forces greater than 6. Round your answers to three decimal places.

What is the lower endpoint and upper endpoint?
Answer #1

sample proportion, = 0.4493
sample size, n = 977
Standard error, SE = sqrt(pcap * (1 - pcap)/n)
SE = sqrt(0.4493 * (1 - 0.4493)/977) = 0.0159

Given CI level is 99%, hence α = 1 - 0.99 = 0.01
α/2 = 0.01/2 = 0.005, Zc = Z(α/2) = 2.58


CI = (pcap - z*SE, pcap + z*SE)
CI = (0.4493 - 2.58 * 0.0159 , 0.4493 + 2.58 * 0.0159)
CI = (0.408 , 0.490)

lower endpoint = 0.408
upper endpoint = 0.490